OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. -4 + 5 + 14 + 23 + ... + 131

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

Do you notice a pattern?

OpenStudy (anonymous):

they go up by 9

OpenStudy (anonymous):

what does the question want you to answer

OpenStudy (anonymous):

to write summation notation for the sequence, filling out \[\sum_{?}^{?} \]

OpenStudy (anonymous):

sorry i can't help you there :(

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

I believe it is\[\sum_{n=0}^{\infty} \] with (-4 + 9n) to the side...maybe

ganeshie8 (ganeshie8):

It is an arithmetic progression with first term = -4 common difference = 9 can you write the general term ?

ganeshie8 (ganeshie8):

Yes, -4 + 9n is the general term !

ganeshie8 (ganeshie8):

How many terms are there in total ?

ganeshie8 (ganeshie8):

-4 + 9n = 131 9n = 135 n = 15 so "n" ranges from 0 to 15, right ?

OpenStudy (anonymous):

i'm not sure how you figure that out? the sequence goes from -4 to 131 i think but can't it keep going to infinity?

ganeshie8 (ganeshie8):

-4 + 5 + 14 + 23 + ... + 131 Notice that the sum ends with the term 131 Its not going to infinity.

OpenStudy (anonymous):

oh ok that makes sense

ganeshie8 (ganeshie8):

-4 + 5 + 14 + 23 + ... + 131 = \(\sum\limits_{n=0}^{15}(-4+9n)\)

ganeshie8 (ganeshie8):

see if that looks good

OpenStudy (anonymous):

if it was going to infinity it would have '...' at the end of the sequence?

ganeshie8 (ganeshie8):

Exactly !

OpenStudy (anonymous):

ok thank you for all of your help! :)